# X-GRM: Large Gaussian Reconstruction Model for Sparse-view X-rays to Computed Tomography

### v1

@@ Abstract
Computed Tomography serves as an indispensable tool in clinical workflows, providing non-invasive visualization of internal anatomical structures. Existing CT reconstruction works are limited to small-capacity model architecture, inflexible volume representation, and small-scale training data. In this paper, we present X-GRM (X-ray Gaussian Reconstruction Model), a large feedforward model for reconstructing 3D CT from sparse-view 2D X-ray projections. X-GRM employs a scalable transformer-based architecture to encode an arbitrary number of sparse X-ray inputs, where tokens from different views are integrated efficiently. Then, tokens are decoded into a new volume representation, named Voxel-based Gaussian Splatting (VoxGS), which enables efficient CT volume extraction and differentiable X-ray rendering. To support the training of X-GRM, we collect ReconX-15K, a large-scale CT reconstruction dataset containing around 15,000 CT/X-ray pairs across diverse organs, including the chest, abdomen, pelvis, and tooth etc. This combination of a high-capacity model, flexible volume representation, and large-scale training data empowers our model to produce high-quality reconstructions from various testing inputs, including in-domain and out-domain X-ray projections. Project Page: https://github.com/CUHK-AIM-Group/X-GRM.
@@ 1 Introduction
Reconstructed from penetrative X-ray projections, Computed Tomography (CT) serves as an indispensable tool in disease diagnosis, treatment planning, and intraoperative guidance x_ray_1 ; x_ray_3 ; x_ray_4 . Traditional reconstruction methods fdk ; analytical_2 ; asd_pocs ; sart typically require several hundred X-ray projections to achieve diagnostically acceptable image quality, exposing patients to excessive radiation and also consuming substantial resources. To address these limitations, our work investigates reconstructing CT volumes from sparse-view X-ray projections.
@@
Existing sparse-view CT reconstruction approaches can be categorized into optimization-based and regression-based methods. Optimization-based methods iteratively refine 3D volumes to match X-ray projections using neural representations zha2022naf ; shen2022nerp ; zha2024r ; cai2024radiative ; sax_nerf or diffusion models chung2023solving ; chung2023decomposed ; lee2023improving . While effective, these approaches typically require minutes to even hours for reconstructing per case, making them impractical for time-sensitive applications. Regression-based methods fbpconvnet leverage neural networks to learn useful patterns for diverse reconstruction tasks-ranging from projection extrapolation anirudh2018lose ; ghani2018deep and slice denoising wang2022dudotrans ; jin2017deep ; freeseed to direct volume regression difnet ; dif_gs ; c2rv . Despite enabling rapid inference, these approaches still face significant challenges: (1) they predominantly employ small model architectures, such as CNNs, which suffer from inherent scaling limitations, (2) they primarily represent CT volumes as 3D voxel, which does not support differentiable X-ray rendering for downstream applications, and (3) they are trained on small-scale, single-organ datasets setio2017validation ; cipriano2022deep ; mccollough2016tu , limiting their applicability to CT volumes with diverse anatomical structures.
@@
Figure 1: Our method achieves state-of-the-art reconstruction quality while maintaining the fastest runtime. (a) Qualitative results: DIF-Gaussian dif_gs exhibits issues with over-smooth results (red boxes) and R2-Gaussian r2_gaussian has noise artifacts (orange boxes) and is time-consuming. In contrast, our method achieves better fidelity in a much shorter time. (b) Performance and runtime comparison: metrics are evaluated on the ReconX-15K test set.
@@
To tackle these limitations, we propose a large feedforward X-ray Gaussian Reconstruction Model (X-GRM), for CT reconstruction from sparse-view X-ray projections. Our innovation is that a transformer-based architecture transformer can excel at CT reconstruction when coupled with training on large-scale datasets. Specifically, we develop a scalable X-ray Reconstruction Transformer that employs an encoder ViT to efficiently tokenize each X-ray projection in parallel, followed by a fusion ViT that integrates global information by performing self-attention across views. Meanwhile, inspired by the efficient rendering of 3D Gaussian Splatting kerbl20233d , we introduce a new volume representation, named Voxel-based Gaussian Splatting (VoxGS), where isometric 3D Gaussians are placed at voxel centers and rendering-related attributes are regressed from tokens. Such design enables efficient CT volume extraction and differentiable X-ray rendering, thereby introducing more constraints from X-ray inputs during training and also supports downstream render-related applications. Finally, to support the training of X-GRM, we introduce a large-scale CT reconstruction dataset ReconX-15K, containing approximately 15,000 volume-projection pairs with diverse anatomical structures. Unlike previous small datasets such as LUNA16 setio2017validation , which contains fewer than 1,000 cases focused on only chest regions, ReconX-15K is 16 times larger and consists of CT volumes with diverse organs, including the chest, abdomen, pelvis, and tooth etc. As shown in Fig. 1, with the proposed scalable model architecture, flexible volume representation, and large-scale data training, X-GRM drastically outperforms the SOTA methods in terms of both the reconstruction quality and inference speed.
@@
  • We introduce X-GRM, a large feed-forward Transformer model that, given an arbitrary number of sparse X-ray projections, can directly predict the CT volume within one second.
@@
  • We collect a large-scale CT reconstruction dataset, ReconX-15K, containing over 16K samples of volume-projection pairs containing diverse anatomical structures.
  • We demonstrate that X-GRM drastically outperforms existing methods in terms of both the reconstruction quality and inference speed. Our codes, pre-trained models, and datasets will be available to the community.
@@ 2.1 Sparse-View CT reconstruction
Current approaches of sparse-view CT reconstruction fall into two fundamental categories. The first encompasses reconstruction through optimization, progressively adjusting the 3D volume to match acquired projections. Conventional approaches sart ; sauer1993local ; sidky2008image formulate this as a maximum a posteriori estimation problem, while contemporary methods incorporate neural representations zha2022naf ; shen2022nerp ; zha2024r ; cai2024radiative ; ddgs ; dgr ; sax_nerf and diffusion modeling chung2023solving ; chung2023decomposed ; lee2023improving . Though these deliver high-quality results, their substantial computational demands render them unsuitable for time-sensitive clinical applications. The second category involves direct regression methodologies using convolutional neural networks ronneberger2015u to learn patterns from extensive datasets. These systems perform reconstruction tasks including projection enhancement anirudh2018lose ; ghani2018deep , slice denoising wang2022dudotrans ; jin2017deep ; freeseed , or volume estimation difnet ; dif_gs ; c2rv . While offering significantly faster processing speeds, these approaches suffer from inherent limitations in neural architecture design, inflexible volume representation, and training on small-scale data. Our proposed X-GRM specifically addresses these fundamental challenges. Concurrent works X-LRM zhang2025x and DeepSparse lin2025deepsparse also explore large foundation models for CT reconstruction, but our work distinguishes itself through novel architecture and 3D Gaussian volume representation, enabling both efficient reconstruction and high-quality X-ray rendering from arbitrary viewpoints—a capability absent in these alternative approaches.
2.2 Feed-forward 3D reconstruction
@@
Unlike methods based on per-sample optimization mildenhall2021nerf ; kerbl20233d , feed-forward reconstruction approaches aim to recover various 3D representations (e.g., mesh wang2018pixel2mesh ; wu2020pq , implicit fields yu2021pixelnerf , 3DGS xu2024grm ; charatan2024pixelsplat ; chen2024mvsplat , and point maps wang2024dust3r ) through a single forward pass from input image sequences. Specifically, in the domain of feed-forward object generation, LRM hong2023lrm and its variants li2023instant3d ; tochilkin2024triposr ; wang2023pf ; wei2024meshlrm ; zhang2024gs ; he2024lucidfusion ; cai2024baking have achieved significant advances in both performance and efficiency by leveraging large-scale datasets deitke2023objaverse ; deitke2023objaversexl and transformer architectures vaswani2017attention . Similarly, in feed-forward scene reconstruction, DUSt3R wang2024dust3r , pixelSplat charatan2024pixelsplat and subsequent variations mast3r ; fast3r ; mvdust3r ; flare ; cut3r ; vggt ; chen2024mvsplat ; noposplat ; wang2024freesplat ; tang2024hisplat have produced impressive results through continuous scaling up of their approaches. However, CT reconstruction remains in its infancy, lacking sufficient generalization capabilities due to the absence of large-scale reconstruction datasets and scalable models. Our work aims to fill this blank.
@@ 3.2 Overview
where we formulate $f_{\bm{\theta}}$ as a feedforward transformer and $\bm{\theta}$ are learnable parameters optimized from a large-scale training dataset. As illustrated in Fig. 2, our method consists of a large X-ray Reconstruction Transformer to enhance the model scalability and a flexible voxel-based Gaussian Splatting that enables efficient CT volume extraction and differentiable rendering. In this section, we first describe the process of encoding X-ray inputs into compact latent tokens, integrated with camera information (Sec. 3.3). Then, we discuss the design of VoxGS and how to regress Gaussian attributes from encoded tokens (Sec. 3.4). Finally, we detailed the model training from large-scale CT collections (Sec. 3.5).
@@ 3.4 Voxel-based Gaussian Splatting
Existing feedforward models dif_gs typically represent CT as 3D voxel, yet it has limitations of supporting differentiable real-time rendering, thus can hardly leverage the X-ray projection constraints during the training period. Benefiting from highly optimized rasterization, 3DGS kerbl20233d enables real-time differentiable rendering. However, directly representing CT with original 3DGS or modified 3DGS for CT r2_gaussian ; ddgs leads to inaccurate volume extraction and instable optimization for feedforward framework (Tab. 6(c)). Therefore, we propose a new CT representation-Voxel-based Gaussian Splatting (VoxGS) to support precise volume extraction and differentiable X-ray rendering.
@@ 3.5 Training objectives and protocols
Table 1: The statistics of ReconX-15K.
@@ 4.1 Experimental setup
Existing CT reconstruction datasets cipriano2022deep ; mccollough2016tu ; setio2017validation exhibit two limitations: restricted data size (<1,000 cases) and single-organ specificity, thereby compromising the potential for developing generalizable and robust models. Addressing this research gap, we establish ReconX-15K, a large and diverse dataset containing 14,972 CT volumes and their corresponding X-ray projections curated from 8 public datasets (detailed in Tab. 1). It spans the anatomical regions most relevant to clinical applications, covering chest, abdomen, pelvis, and dental structures. To facilitate model development and evaluation, we partition ReconX-15K into three subsets designated as 13,612, 680, and 680, serving as training, validation, and testing sets respectively. Detailed split of each subset is describe in Sec. A.2.
To standardize CT volumes, we resample anatomical regions to consistent dimensions: chest, abdomen, and pelvis volumes to $50^{3}\text{cm}^{3}$ with $256^{3}$ voxels; dental volumes to $40^{3}\text{cm}^{3}$ with $256^{3}$ voxels. Radiodensity values are normalized from Hounsfield units ([-1000, 1000] for chest, abdomen, pelvis; [-1000, 3000] for dental structures) to the range [0,1], ensuring coverage of primary organs of interest. As most public datasets provide only CT volumes, we generated multi-view projections using the TIGRE biguri2016tigre toolbox. For each volume, we rendered 50 X-ray projections at $256^{2}$ resolution, evenly distributed across the full $[0^{\circ},360^{\circ}]$ range. To enhance realism, we incorporated Gaussian and Poisson noise to simulate Compton scattering effects.
@@
During model training, we employ a variable-view learning strategy with the number of views randomly selected from $\{6,8,10\}$, and views are evenly sampled from the 50 available projections. For quantitative evaluation, we utilize both PSNR (calculated directly in 3D volumetric space) and SSIM ssim (averaged across 2D slice comparisons) to comprehensively assess reconstruction quality through signal fidelity and structural similarities.
Table 2: Quantitative comparison with traditional and feedforward methods. We evaluate all methods using the full ReconX-15K test set (2563 resolution) on one A100 GPU. All feedforward models are trained on the same training set. Bold: best results; underlined: second-best results.
@@ 4.2 Assessing CT reconstruction
We evaluate our X-GRM against existing methods with two settings. Setting 1: evaluating on the full test set (680 samples) of ReconX-15K. This setting includes traditional and feedforward methods, as they can efficiently handle massive X-ray inputs. To ensure fair comparison, we re-train the compared feedforward models on the same training set. Setting 2: evaluating on a sampled set (40 samples). This setting includes self-supervised approaches, as they typically require substantial optimization time. To cover diverse organs, we evenly sample five volumes from the test set of eight datasets.
@@ 4.3 Assessing cross-dataset generalization
Due to the use of scalable architecture and large-scale training, X-GRM is inherently superior in generalizing to out-of-distribution inputs. To demonstrate this advantage, we conduct two cross-dataset evaluations. In particular, we use the models trained on ReconX-15K, and directly test them on unseen PENGWIN liu2023pelvic (pelvis) and FUMPE masoudi2018 (chest) datasets. As evident in Tab. 4, our X-GRM significantly outperforms other feedforward models, including FreeSeed, DIF-Net, and DIF-Gaussian. Compared to self-supervised models that uses per-sample optimization, our method also shows comparable or better performance, while delivering much faster inference speed (around 500× faster compared to R2-Gaussian).

### v2

@@ Abstract
Computed Tomography serves as an indispensable tool in clinical workflows, providing non-invasive visualization of internal anatomical structures. Existing CT reconstruction works are limited to small-capacity model architecture and inflexible volume representation. In this work, we present X-GRM (X-ray Gaussian Reconstruction Model), a large feedforward model for reconstructing 3D CT volumes from sparse-view 2D X-ray projections. X-GRM employs a scalable transformer-based architecture to encode sparse-view X-ray inputs, where tokens from different views are integrated efficiently. Then, these tokens are decoded into a novel volume representation, named Voxel-based Gaussian Splatting (VoxGS), which enables efficient CT volume extraction and differentiable X-ray rendering. This combination of a high-capacity model and flexible volume representation, empowers our model to produce high-quality reconstructions from various testing inputs, including in-domain and out-domain X-ray projections. Our codes are available at: https://github.com/CUHK-AIM-Group/X-GRM.
@@ 1 Introduction
Computed Tomography (CT) serves as an indispensable tool in disease diagnosis, treatment planning, and intraoperative guidance x_ray_1 ; x_ray_3 ; x_ray_4 . Traditional reconstruction methods fdk ; analytical_2 ; asd_pocs ; sart typically require several hundred X-ray projections to achieve diagnostically acceptable image quality, exposing patients to excessive radiation and also consuming substantial resources. To address these limitations, our work investigates reconstructing CT volumes from sparse-view X-ray projections.
@@
Existing sparse-view CT reconstruction approaches can be categorized into optimization-based and regression-based methods. Optimization-based methods iteratively refine 3D volumes to match X-ray projections using neural representations zha2022naf ; shen2022nerp ; zha2024r ; cai2024radiative ; sax_nerf or diffusion models chung2023solving ; chung2023decomposed ; lee2023improving . While effective, these approaches typically require minutes to even hours for reconstructing per case, making them impractical for time-sensitive applications. Regression-based methods fbpconvnet leverage neural networks to learn useful patterns for diverse reconstruction tasks-ranging from projection extrapolation anirudh2018lose ; ghani2018deep and slice denoising wang2022dudotrans ; jin2017deep ; freeseed to direct volume regression difnet ; dif_gs ; c2rv . Despite enabling rapid inference, these approaches still face significant challenges: (1) they primarily employ small model architectures, like convolutional neural networks (CNNs), suffering from inherent scaling limitations, and (2) they primarily represent CT volumes as 3D voxels, which do not support differentiable X-ray rendering for incorporating X-ray constraints and render-related applications.
Very recently, X-LRM zhang2025x firstly proposes to use scalable transformer architectures trained on extensive datasets, while DeepSparse lin2025deepsparse similarly explores large-scale models for CT reconstruction. Both approaches achieve impressive reconstruction quality with computational efficiency. However, these methods rely on modified neural radiance fields zhang2025x or discrete 3D voxel representations lin2025deepsparse , which lack support for differentiable X-ray rendering—limiting their ability to incorporate direct X-ray constraints and rendering-based applications.
@@
Figure 1: Our method achieves state-of-the-art reconstruction quality while maintaining the fastest runtime. (a) Qualitative results: DIF-Gaussian dif_gs exhibits issues with over-smooth results (red boxes) and R2-Gaussian r2_gaussian has noise artifacts (orange boxes) and is time-consuming. In contrast, our method achieves better fidelity in a much shorter time. (b) Performance and runtime comparison: metrics are evaluated on the test set of collected large-scale dataset.
@@
To tackle these limitations, we propose a large feedforward X-ray Gaussian Reconstruction Model (X-GRM), for CT reconstruction from sparse-view X-ray projections. Our innovation is that a transformer-based architecture transformer , when coupled with flexible volume representations, can excel at CT reconstruction. Specifically, we develop a scalable X-ray Reconstruction Transformer that employs an encoder ViT to efficiently tokenize each X-ray projection in parallel, followed by a fusion ViT that integrates global information by performing self-attention across views. Secondly, inspired by the efficient rendering of 3D Gaussian Splatting kerbl20233d , we introduce a new volume representation, named Voxel-based Gaussian Splatting (VoxGS), where isometric 3D Gaussians are placed at voxel centers and rendering-related attributes are regressed from tokens. Such design enables efficient CT volume extraction and differentiable X-ray rendering, thereby introducing more constraints from X-ray inputs during training and also supports downstream render-related applications. As shown in Fig. 1, with the proposed scalable architecture and flexible volume representation, X-GRM drastically outperforms the SOTA methods in terms of both the reconstruction quality and inference speed.
@@
  • We introduce X-GRM, a large feed-forward Transformer model that, given sparse X-ray projection inputs, can directly predict the CT volume within one second.
@@
  • We demonstrate that X-GRM drastically outperforms existing methods in both the reconstruction quality and inference speed. Our codes and pre-trained models will be available.
@@ 2.1 Sparse-view CT reconstruction
Sparse-view CT reconstruction tackles the complex task of creating complete 3D volumetric data from a limited number of X-ray projections. Traditional methods sart ; sauer1993local ; sidky2008image frame this as maximum a posteriori estimation with statistical priors, but often produce poor quality reconstructions when very few views are available. Recent advances have employed neural radiance fields zha2022naf ; shen2022nerp ; sax_nerf or 3D Gaussian primitives cai2024radiative ; ddgs ; dgr ; zha2024r as volume representations, which are gradually refined until their rendered X-rays match the input projections. Other approaches explore diffusion models as priors chung2023solving ; chung2023decomposed ; lee2023improving . While these methods deliver better results than traditional algorithms, they require extensive optimization time for each case, making them impractical for urgent clinical needs.
Feed-forward models offer a promising solution by directly mapping sparse projections to complete volumes with much faster inference times. These approaches include sinogram enhancement anirudh2018lose ; ghani2018deep , cross-sectional image denoising wang2022dudotrans ; jin2017deep ; freeseed , and direct volume synthesis difnet ; dif_gs ; c2rv . Though significantly faster, these methods are limited by their architectural design and rigid volume representations. Recent works like X-LRM zhang2025x and DeepSparse lin2025deepsparse explore large foundation models for reconstruction, while our approach stands out through innovative architecture and 3D Gaussian volume representation—enabling both efficient reconstruction and high-quality X-ray rendering from any viewpoint, a capability missing in other approaches.
2.2 Feed-forward 3D object/scene reconstruction
@@
Feed-forward reconstruction approaches aim to recover various 3D representations (e.g., mesh wang2018pixel2mesh ; wu2020pq , implicit fields yu2021pixelnerf , 3DGS xu2024grm ; charatan2024pixelsplat ; chen2024mvsplat , and point maps wang2024dust3r ) through a single forward pass from input images. Specifically, in the domain of feed-forward object generation, LRM hong2023lrm and its variants li2023instant3d ; tochilkin2024triposr ; wang2023pf ; wei2024meshlrm ; zhang2024gs ; he2024lucidfusion ; cai2024baking have achieved significant advances in both performance and efficiency by leveraging large-scale datasets deitke2023objaverse ; deitke2023objaversexl and transformer architectures vaswani2017attention . Similarly, in feed-forward scene reconstruction, DUSt3R wang2024dust3r , pixelSplat charatan2024pixelsplat and subsequent variations mast3r ; fast3r ; mvdust3r ; flare ; cut3r ; vggt ; chen2024mvsplat ; noposplat ; wang2024freesplat ; tang2024hisplat have produced impressive results through continuous scaling up of their approaches. However, CT reconstruction remains in its infancy, lacking sufficient generalization capabilities due to the use of small-scale model architectures and inflexible volume representations. In this work, we aim to tackle these limitations.
@@ 3.2 Overview
where we formulate $f_{\bm{\theta}}$ as a feedforward transformer and $\bm{\theta}$ are learnable parameters optimized from a large-scale training dataset. As illustrated in Fig. 2, our method consists of a large X-ray Reconstruction Transformer to enhance the model scalability and a flexible voxel-based Gaussian Splatting that enables efficient CT volume extraction and differentiable rendering. In this section, we first describe the process of encoding X-ray inputs into compact latent tokens, integrated with camera information (Sec. 3.3). Then, we discuss the design of VoxGS and how to regress Gaussian attributes from encoded tokens (Sec. 3.4). Finally, we describe the model training details (Sec. 3.5).
@@ 3.4 Voxel-based Gaussian Splatting
Existing feedforward models dif_gs typically represent CT as 3D voxels, yet it has limitations of supporting differentiable real-time rendering, thus can hardly leverage the X-ray projection constraints during the training period. Benefiting from highly optimized rasterization, 3DGS kerbl20233d enables real-time differentiable rendering. However, directly representing CT with original 3DGS or modified 3DGS for CT r2_gaussian ; ddgs leads to inaccurate volume extraction and instable optimization for feedforward framework (Tab. 6(c)). Therefore, we propose a new CT representation-Voxel-based Gaussian Splatting (VoxGS) to support precise volume extraction and differentiable X-ray rendering.
@@ 3.5 Training objectives and protocols
Table 1: The statistics of collected datasets.
@@ 4.1 Experimental setup
To support the training of our scalable transformer with diverse anatomical data, we assemble an extensive dataset following X-LRM’s zhang2025x selection, combining 14,972 CT volumes and their corresponding X-ray projections from 8 public datasets (detailed in Tab.1). This comprehensive repository spans the most clinically relevant body regions, including chest, abdomen, pelvis, and dental structures. For systematic model development and evaluation, we divided the collection into three distinct subsets: 13,612 volumes for training, 680 for validation, and 680 for testing. Complete specifications of these dataset partitions are thoroughly documented in Sec.A.2.
For CT standardization, we meticulously normalize anatomical regions: chest, abdomen, and pelvic volumes are resampled to $50^{3}\text{cm}^{3}$ at $256^{3}$ voxel resolution, while dental acquisitions are calibrated to $40^{3}\text{cm}^{3}$ with the same resolution. We also transform radio-density measurements from native Hounsfield units (chest, abdomen, pelvis: [-1000, 1000]; dental: [-1000, 3000]) to a unified [0,1] scale, preserving crucial diagnostic information across key structures. To render X-ray projections from CT volumes in each dataset, we employ the TIGRE biguri2016tigre toolbox to generate synthetic radiographic projections—producing 50 high-definition views ($256^{2}$ resolution) distributed uniformly throughout the complete angular spectrum ($[0^{\circ},360^{\circ}]$). Clinical authenticity was enhanced through noise integration, combining Gaussian and Poisson distributions to accurately replicate quantum photon effects and Compton scatter phenomena.
@@
During model training, we employ a variable-view learning strategy with the number of views randomly selected from ${6,8,10}$, with these perspectives evenly sampled from the 50 available projections. For quantitative evaluation, we utilize complementary metrics: PSNR calculated directly within the 3D volumetric space provides assessment of signal fidelity, while SSIM ssim averaged across multiple 2D slice comparisons offers insight into structural preservation, together providing a comprehensive measurement of reconstruction quality.
Table 2: Quantitative comparison with traditional and feedforward methods. We evaluate all methods using the full test set (2563 resolution) on one A100 GPU. All feedforward models are trained on the same training set. Bold: best results; underlined: second-best results.
@@ 4.2 Assessing CT reconstruction
Following the setup of X-LRM zhang2025x , we evaluate our X-GRM against existing CT reconstruction methods with two settings. Setting 1: evaluating on the full test set (680 samples) of the collected dataset. This setting includes traditional and feedforward methods, as they can efficiently handle massive X-ray inputs. To ensure fair comparison, we re-train the compared feedforward models on the same training set. Setting 2: evaluating on a sampled set (40 samples). This setting includes self-supervised approaches, as they typically require substantial optimization time. To cover diverse organs, we evenly sample five volumes from the test set of eight datasets.
@@ 4.3 Assessing cross-dataset generalization
Due to the use of scalable architecture and large-scale training, X-GRM is inherently superior in generalizing to out-of-distribution inputs. To demonstrate this advantage, we conduct two cross-dataset evaluations. In particular, we use the models trained on the collected dataset, and directly test them on unseen PENGWIN liu2023pelvic (pelvis) and FUMPE masoudi2018 (chest) datasets. As evident in Tab. 4, our X-GRM significantly outperforms other feedforward models, including FreeSeed, DIF-Net, and DIF-Gaussian. Compared to self-supervised models that uses per-sample optimization, our method also shows comparable or better performance, while delivering much faster inference speed (around 500× faster compared to R2-Gaussian).